# From System 1 to System 2: A Survey of Reasoning Large Language Models

### v1

@@ 2.1 Foundational LLMs
By 2023-2024, models such as GPT-4 [62], LLaMA [25], and LLaVA [27] demonstrated advanced capabilities in reasoning, contextual understanding, and multimodal reasoning, processing both text and images. The evolution of foundational LLMs has revolutionized AI, enabling more sophisticated applications in language comprehension, problem-solving, and human-machine collaboration.
Summary: The development of foundational LLMs has progressed from pretrained transformers like BERT to multimodal models such as GPT-4, enhancing language understanding, text generation, and image processing. This advancement has led to significant breakthroughs in AI, improving language comprehension, problem-solving, and human-computer interaction. Building on deep learning advancements [63, 64, 65, 66], foundational LLMs can learn extensive world knowledge and semantic relationships from vast textual or multimodal data. This enables them to exhibit emergent capabilities such as In-Context Learning (ICL) [67], prompt engineering [68], and Chain-of-Thought (CoT) reasoning [2], significantly enhancing their adaptability and creative problem-solving abilities.
@@ 3.1.1 Output Behaviour Perspective
Longer Inference Length & Time: Recent research [49, 50, 51, 52] indicates that reasoning LLMs often generate outputs exceeding 2000 tokens to tackle complex problems in coding and mathematics. However, this extended output length can sometimes lead to overthinking, where the model spends excessive time on a problem without necessarily improving the solution. Studies [49] highlight that while autoregressive generation and Classic CoT can effectively solve simpler problems, they struggle with more complex tasks. Research [95, 96] shows that in multimodal domains, many problems demand careful observation, comparison, and deliberation. Additionally, Search-o1 [97] suggests that slow-thinking mechanisms are particularly beneficial in areas requiring external knowledge or where potential knowledge conflicts arise. In medical scenarios, complex problems, such as those requiring test-time scaling techniques, demonstrate significant improvements [52].
@@ 3.2 Core Method
TABLE I: Summary of Structure Search method.
@@
Correction and Exploration RethinkMCTS [111], MCTSr [112] Actions emphasize revisiting, refining, or backtracking to improve previous reasoning steps.
@@ 3.2.1 Structure Search
Reasoning LLMs aim to achieve high accuracy and depth in solving complex problems by emulating the deliberate and methodical nature of human reasoning. However, despite recent advancements, current foundational LLMs face inherent limitations when addressing intricate reasoning tasks. These limitations arise from their lack of an internal world model to simulate environmental states, their inability to predict the long-term outcomes of reasoning paths, and their failure to iteratively refine reasoning steps based on future states or rewards [8]. As a result, these shortcomings hinder foundational LLMs from effectively balancing exploration and exploitation in vast reasoning spaces, creating challenges in tasks that require multi-step reasoning, such as complex mathematics, logical inference, or strategic decision-making [119].
@@
Additionally, as shown in Table I, we classify the reward design into five categories:
@@ 3.2.2 Reward Modeling
Stage Methods Data Source Model Refinement Applications
@@ 3.2.3 Self Improvement
Reasoning LLMs exemplify a progression from weak to strong supervision, while traditional CoT fine-tuning faces challenges in scaling effectively. Self improvement, using the model’s exploration capabilities for self-supervision, gradually enhances LLMs performance in tasks such as translation [148], mathematical reasoning [144, 149], and multimodal perception [152]. This approach fosters exploration and application within reasoning LLMs [165, 147]. A summary of Self Improvement method is presented in Table III.
@@ 3.2.4 Macro Action
TABLE IV: Summary of Macro Action method.
Methods Usage Action Attribute Representative Action
@@
Self-Check[166] Verification Human-Designed 4 ICL TT Target Extraction, Information Collection, Step Regeneration, Result Comparsion
LeMa[167] Synthetic Data Human-Designed 3 ICL & SFT TT Incorrect Step Recognition, Explanation, Correct Solution:
REFINER[168] Verification/Exploration Human-Designed 2 ICL & SFT TT Critic, Generate
HiICL-MCTS[169] Exploration Human-Designed 5 ICL TT System Analysis, One-Step Thought, Divide and Conquer, …, Self-Reflection and Refinement
SUPERCORRECT[170] Distill In-Context Learning Dynamic SFT & RL TT
ReasonFlux[171] Synthetic Data/Exploration Human-Designed $\sim$500 ICL & SFT & RL TT
rStar[172] Exploration Human-Designed 5 ICL & RL TT One-step thought, Propose Next Sub-question & Answer, …, Rephrase question
LLaMA-Berry[173] Exploration Human-Designed 2 ICL & RL TT Reflection, Error Re-correction
Huatuo-o1[94] Synthetic Data Human-Designed 4 ICL & SFT TT Backtracking, Exploring New Paths, Verification, Correction
Marco-o1[93] Verification Human-Designed 1 ICL & SFT TT Reflection
BoT[174] Exploration In-Context Learning Dynamic ICL TT Solving Quadratic Equation, Array Sorting, …, Search Algorithms)
rStar-Math[147] Exploration In-Context Learning $1$ ICL & RL TT Python comment
Mulberry[175] Synthetic Data In-Context Learning 1 ICL & SFT TT Reflection
LLaVA-CoT[176] Synthetic Data/Exploration Human-Designed 4 SFT II TT Summary, Caption, Reasoning, Conclusion
LLaMAV-o1[177] Verification/Exploration Human-Designed 4173 Curriculum Learning II TT Detailed Caption Generation, Logical Reasoning, … Final Answer Generation
AtomThink[178] Synthetic Data/Exploration In-Context Learning $>$100 SFT & RL II TT Variable Definition, Calculations, Graphs Analysis , …, Verification
RedStar[54] Distill Human-Designed 2 SFT II TT Wait, Alternately
Auto-CoT[179] Exploration In-Context Learning 2 ICL TT Question clustering, Demonstration Sampling
PoT[180] Verification In-Context Learning 1 ICL TT Code language conversion
PAL[181] Verification In-Context Learning 1 ICL TT Code language conversion
Decomposed Prompt[182] Exploration Human-Designed 3 ICL TT Peoblem Split, Subproblem Solving, Answer Merge
Least-to-Most[183] Exploration Human-Designed 2 ICL TT Problem Decomposition, Subproblem Solving
@@
Macro actions also play a crucial role in enhancing Self-improvement frameworks. rStar-Math [147] utilizes high-level deliberate search through Code-augmented CoT, generating diverse and reliable solutions while achieving proactive search capabilities. Satori [186] integrates CoT with RL, incorporating “$<$reflect$>$”-style macro actions to diversify exploration and alleviate policy saturation in online RL environments. Huatuo-o1 [94] combines hierarchical planning with domain-specific knowledge bases to improve medical reasoning. Additionally, ReasonFlux [171] dynamically reconfigures reasoning templates (e.g., breaking down calculus problems into symbolic and numeric phases) to align with the problem structure.
@@ 3.2.5 Reinforcement Fine-Tuning
Reinforcement Fine-Tuning (RFT) [187] is an innovative technique recently introduced by OpenAI, designed to enable developers and engineers to fine-tune existing models for specific domains or complex tasks. Unlike general SFT, RFT focuses on optimizing the model’s reasoning process by using a reward mechanism to guide the model’s evolution, thereby enhancing its reasoning capabilities and accuracy. The core of RFT lies in improving the model’s performance in a specific domain with minimal high-quality training data [188], an appropriate reward model [189], and a stable optimization process [190]. A summary of RFT method is presented in Table V.
@@
TABLE V: Summary of RFT method.
@@
Foundational LLMs Modality Reward Type Algorithm Learning Incentivize Sample
@@ 3.3 Evolutionary of Reasoning LLMs
TABLE VI: Statistics of benchmarks for reasoning LLMs.
Domain Benchmark Venue Language Size Level
Math AIME 2024 [226] - English 30 Competition
MATH-500 [37] ICLR 2024 English 500 Competition
AMC 2023 [227] - English 30 Competition
Olympiad Bench [228] ACL 2024 English/Chinese 8,476 Competition
Code Codeforces - English - Expert
SWE-bench [229] ICLR 2024 English 2,294 Expert
LiveCodeBench [230] ArXiv 2024 English - Expert
Science GPQA Diamond [231] COLM 2024 English 448 University
MMLU-Pro [232] NeurIPS 2024 English 12,032 Hybrid
Agent WebShop [233] NeurIPS 2022 English 1,600 Hybrid
WebArena [234] ICLR 2024 English 812 Hybrid
SciWorld [235] EMNLP 2022 English 7,200 Hybrid
TextCraft [236] NAACL 2024 English 200 Hybrid
Medicine JAMA Clinical Challenge [237] NAACL 2025 English 1,524 Expert
Medbullets [237] NAACL 2025 English 308 Expert
MedQA [238] ArXiv 2020 English/Chinese 61,097 Expert
Multimodality MMMU [239] CVPR 2024 English 11,500 Hybrid
MathVista [240] ICLR 2024 English 6,141 Middle School
MathVision [241] NeurIPS 2024 English 3,040 Middle/High School
CMMaTH [242] COLING 2025 English/Chinese 23,856 Middle/High School
PGPS9K [243] IJCAI 2023 English 9,023 Middle School
@@ 4.2 Evaluation Metrics
4.2.1 Task Types
4.2.2 Technical Proposals
4.2.3 Reasoning Paradigms
4.2.4 Summary
4.3 Performance Comparison
In this section, we compare the performance of different reasoning LLMs and their corresponding foundational LLMs on plain text benchmarks, such as math and code problems, as well as on multimodal benchmarks. The comprehensive real-time leaderboard is available on this website.
4.3.1 Performance on Plain Text Benchmarks
4.3.2 Performance on Multimodal Benchmarks
@@ 5.2 Collaborative Slow & Fast-thinking Systems
A key challenge in reasoning LLMs is the loss of fast-thinking capabilities, which results in inefficiencies when simple tasks require unnecessary deep reasoning. Unlike humans, who fluidly switch between fast (System 1) and slow (System 2) thinking, current reasoning LLMs struggle to maintain this balance. While reasoning LLMs ensure deliberate and thorough reasoning, fast-thinking systems rely on prior knowledge for quick responses. Despite efforts such as the System 1-2 switcher [95], speculative decoding [258, 259, 260], and interactive continual learning [261], integrating both modes of thinking remains challenging. This often leads to inefficiencies in domain-specific tasks and underutilized strengths in more complex scenarios.
@@ 5.6 Multimodal Reasoning LLMs
Extending slow-thinking reasoning capabilities from text-based domains to multimodal contexts remains a significant challenge, especially in tasks requiring fine-grained perception [96]. While approaches like Virgo [269] have attempted to distill text-based slow-thinking reasoning into multimodal LLMs, their performance improvements in tasks such as MathVision [241], which demand detailed visual understanding, have been marginal.
@@ 5.7 Safe Reasoning LLMs
The rapid development of reasoning LLMs like OpenAI-o1 [29] and DeepSeek-R1 [31] has led to the rise of superintelligent models capable of continuous self-evolution. However, this progress brings challenges in safety and control. RL, a key training method, introduces risks such as reward hacking, generalization failures, and language mixing, which can lead to harmful outcomes. Ensuring the safety of such systems like DeepSeek-R1 is urgent. While RL enhances reasoning, its uncontrollable nature raises concerns about safely guiding these models. SFT addresses some issues but is not a complete solution. A hybrid approach combining RL and SFT is needed to reduce harmful outputs while maintaining model effectiveness [270].

### v2

@@ 2.1 Foundational LLMs
By 2023-2024, models such as GPT-4/4o [62, 16], LLaMA [25], and LLaVA [27] demonstrated advanced capabilities in reasoning, contextual understanding, and multimodal reasoning, processing both text and images [63, 64, 65]. DeepSeek-V3 [17], featuring a 671B Mixture-of-Expert architecture [66, 67, 68], outperforms several other LLMs on key benchmarks while offering significant improvements in efficiency and processing speed. The evolution of foundational LLMs has revolutionized AI, enabling more sophisticated applications in language comprehension, problem-solving, and human-machine collaboration.
Summary: The development of foundational LLMs has progressed from pretrained transformers like BERT to multimodal models such as GPT-4, enhancing language understanding, text generation, and image processing. This advancement has led to significant breakthroughs in AI, improving language comprehension, problem-solving, and human-computer interaction. Building on deep learning advancements [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 18, 83], foundational LLMs can learn extensive world knowledge and semantic relationships from vast textual or multimodal data. This enables them to exhibit emergent capabilities such as In-Context Learning (ICL) [84, 85], prompt engineering [86, 87], and Chain-of-Thought (CoT) reasoning [2], significantly enhancing their adaptability and creative problem-solving abilities.
@@ 3.1.1 Output Behaviour Perspective
Longer Inference Length & Time: Recent research [49, 50, 51, 52, 114] indicates that reasoning LLMs often generate outputs exceeding 2000 tokens to tackle complex problems in coding and mathematics. However, this extended output length can sometimes lead to overthinking, where the model spends excessive time on a problem without necessarily improving the solution. Studies [49] highlight that while autoregressive generation and Classic CoT can effectively solve simpler problems, they struggle with more complex tasks. Research [115, 116] shows that in multimodal domains, many problems demand careful observation, comparison, and deliberation. Additionally, Search-o1 [117] suggests that slow-thinking mechanisms are particularly beneficial in areas requiring external knowledge or where potential knowledge conflicts arise. In medical scenarios, complex problems, such as those requiring test-time scaling techniques, demonstrate significant improvements [52].
@@ 3.2 Core Method
TABLE I: Summary of Structure Search method based on the definition of actions and rewards.
@@
Correction and Exploration RethinkMCTS [131], MCTSr [132] Actions focus on revisiting and refining previous steps.
@@ 3.2.1 Structure Search
Reasoning LLMs aim to achieve high accuracy and depth in solving complex problems by emulating the deliberate nature of human reasoning. However, despite recent advancements, current foundational LLMs face inherent limitations when addressing intricate reasoning tasks. These limitations arise from their lack of an internal world model to simulate environmental states, their inability to predict the long-term outcomes of reasoning paths, and their failure to iteratively refine reasoning steps based on future states or rewards [8]. As a result, these shortcomings hinder foundational LLMs from effectively balancing exploration and exploitation in vast reasoning spaces, creating challenges in tasks that require multi-step reasoning, such as complex mathematics, logical inference, or strategic decision-making [139].
@@
Additionally, as illustrated in Table I, we classify the reward design into five categories:
@@ 3.2.2 Reward Modeling
Stage Methods Data Source Model Refinement Application
@@ 3.2.3 Self Improvement
Reasoning LLMs exemplify a progression from weak to strong supervision, while traditional CoT fine-tuning faces challenges in scaling effectively. Self improvement, using the model’s exploration capabilities for self-supervision, gradually enhances LLMs performance in tasks such as translation [167], mathematics [163, 168], and multimodal perception [171]. This approach fosters exploration and application within reasoning LLMs [184, 185, 186, 166]. A summary of Self Improvement method is presented in Table III.
@@ 3.2.4 Macro Action
TABLE IV: Summary of Macro Action method. Tag: II = Image, TT = Text, VV = Video. Action Category: AD: Analysis and Decomposition, IPR: Information Processing and Reasoning, VC: Verification and Correction, GO: Generation and Optimization, EB: Exploration and Backtracking.
Methods Usage Action Attribute Main Action Category
@@
Self-Check[187] Verification Human-Designed 4 ICL TT AD, VC
LeMa[188] Synthetic Data Human-Designed 3 ICL & SFT TT VC, IPR
REFINER[189] Verification/Exploration Human-Designed 2 ICL & SFT TT VC, AD
HiICL-MCTS[190] Exploration Human-Designed 5 ICL TT VC, EB, AD
SUPERCORRECT[191] Distill In-Context Learning Dynamic SFT & RL TT AD, IPR
ReasonFlux[192] Synthetic Data/Exploration Human-Designed $\sim$500 ICL & SFT & RL TT AD, IPR
rStar[193] Exploration Human-Designed 5 ICL & RL TT VC, GO, EB
LLaMA-Berry[194] Exploration Human-Designed 2 ICL & RL TT VC, EB
Huatuo-o1[113] Synthetic Data Human-Designed 4 ICL & SFT TT VC
Marco-o1[112] Verification Human-Designed 1 ICL & SFT TT VC
BoT[195] Exploration In-Context Learning Dynamic ICL TT AD, IPR
rStar-Math[166] Exploration In-Context Learning $1$ ICL & RL TT AD, IPR
Mulberry[196] Synthetic Data In-Context Learning 1 ICL & SFT TT VC, EB
LLaVA-CoT[197] Synthetic Data/Exploration Human-Designed 4 SFT II TT AD, IPR
LLaMAV-o1[198] Verification/Exploration Human-Designed 4173 Curriculum Learning II TT AD, IPR
AtomThink[199] Synthetic Data/Exploration In-Context Learning $>$100 SFT & RL II TT AD, IPR, EB
RedStar[54] Distill Human-Designed 2 SFT II TT AD, VC
Auto-CoT [13] Exploration In-Context Learning 2 ICL TT AD, IPR, GO
PoT[200] Verification In-Context Learning 1 ICL TT AD, IPR, GO
PAL[201] Verification In-Context Learning 1 ICL TT AD, IPR, GO
Decomposed Prompt[202] Exploration Human-Designed 3 ICL TT AD, IPR
Least-to-Most[203] Exploration Human-Designed 2 ICL TT AD, IPR
@@
Macro actions also play a crucial role in enhancing self improvement frameworks. rStar-Math [166] utilizes high-level deliberate search through Code-augmented CoT, generating diverse and reliable solutions while achieving proactive search capabilities. Satori [206] integrates CoT with RL, incorporating “$<$reflect$>$”-style macro actions to diversify exploration and alleviate policy saturation in online RL environments. Huatuo-o1 [113] combines hierarchical planning with domain-specific knowledge bases to improve medical reasoning. Additionally, ReasonFlux [192] dynamically reconfigures reasoning templates (e.g., breaking down calculus problems into symbolic and numeric phases) to align with the problem structure.
@@ 3.2.5 Reinforcement Fine-Tuning
Reinforcement Fine-Tuning (RFT) [207] is an innovative technique recently introduced by OpenAI, designed to enable developers and engineers to fine-tune existing models for specific domains or complex tasks. Unlike general SFT, RFT focuses on optimizing the model’s reasoning process by using a reward mechanism to guide the model’s evolution, thereby enhancing its reasoning capabilities and accuracy. The core of RFT lies in improving the model’s performance in a specific domain with minimal high-quality training data [208], an appropriate reward model [209], and a stable optimization process in long-context [210, 211, 212, 213]. A summary of RFT method is presented in Table V.
@@
TABLE V: Summary of RFT method. Tag: II = Image, TT = Text, VV = Video.
@@
Foundational LLMs Feedback Modality Reward Type Algorithm Learning Incentivize Sample
@@ 3.3 Evolutionary of Reasoning LLMs
TABLE VI: Statistics of benchmarks for reasoning LLMs.
Domain Benchmark Question Type Venue Language Size Level
Math AIME 2024 [246] Open-End - English 30 Competition
MATH-500 [37] Open-End ICLR 2024 English 500 Competition
AMC 2023 [247] Open-End English 30 Competition
Olympiad Bench [248] Open-End ACL 2024 English/Chinese 8,476 Competition
Code Codeforces Open-End - English - Expert
SWE-bench [249] Open-End ICLR 2024 English 2,294 Expert
LiveCodeBench [250] Open-End ArXiv 2024 English - Expert
Science GPQA Diamond [251] Choice COLM 2024 English 448 University
MR-Ben [252] Hybrid NeurIPS 2024 English 5,975 Hybrid
MMLU-Pro [253] Choice NeurIPS 2024 English 12,032 Hybrid
Agent WebShop [254] Open-End NeurIPS 2022 English 1,600 Hybrid
WebArena [255] Open-End ICLR 2024 English 812 Hybrid
SciWorld [256] Open-End EMNLP 2022 English 7,200 Hybrid
TextCraft [257] Open-End NAACL 2024 English 200 Hybrid
Medicine JAMA Clinical [258] Choice NAACL 2025 English 1,524 Expert
Medbullets [258] Choice NAACL 2025 English 308 Expert
MedQA [259] Choice ArXiv 2020 English/Chinese 61,097 Expert
Multimodality MMMU [260] Hybrid CVPR 2024 English 11,500 Hybrid
MathVista [261] Hybrid ICLR 2024 English 6,141 Middle School
MathVision [262] Hybrid NeurIPS 2024 English 3,040 Middle/High School
CMMaTH [263] Hybrid COLING 2025 English/Chinese 23,856 Middle/High School
PGPS9K [264] Hybrid IJCAI 2023 English 9,023 Middle School
@@ 4.2 Evaluation Metrics
4.2.1 Task Types
4.2.2 Technical Proposals
4.2.3 Reasoning Paradigms
4.2.4 Summary
4.3 Performance Comparison
In this section, we compare the performance of different reasoning LLMs and their corresponding foundational LLMs on plain text benchmarks, such as math and code problems, as well as on multimodal benchmarks. The comprehensive real-time leaderboard is available on Arena.
4.3.1 Performance on Plain Text Benchmarks
4.3.2 Performance on Multimodal Benchmarks
@@ 5.2 Collaborative Slow & Fast-thinking Systems
A key challenge in reasoning LLMs is the loss of fast-thinking capabilities, which results in inefficiencies when simple tasks require unnecessary deep reasoning. Unlike humans, who fluidly switch between fast (System 1) and slow (System 2) thinking, current reasoning LLMs struggle to maintain this balance. While reasoning LLMs ensure deliberate and thorough reasoning, fast-thinking systems rely on prior knowledge for quick responses. Despite efforts such as the System 1-2 switcher [115], speculative decoding [282, 283, 284], and interactive continual learning [285, 286, 287], integrating both modes of thinking remains challenging. This often leads to inefficiencies in domain-specific tasks and underutilized strengths in more complex scenarios.
@@ 5.6 Multimodal Reasoning LLMs
Extending slow-thinking reasoning capabilities from text-based domains to multimodal contexts remains a significant challenge, especially in tasks requiring fine-grained perception [116, 295, 296]. While approaches like Virgo [297] have attempted to distill text-based slow-thinking reasoning into multimodal LLMs, their performance improvements in tasks such as MathVision [262], which demand detailed visual understanding, have been marginal.
@@ 5.7 Safe Reasoning LLMs
The rapid development of reasoning LLMs like OpenAI-o1 [29] and DeepSeek-R1 [31] has led to the rise of superintelligent models capable of continuous self-evolution. However, this progress brings challenges in safety and control [298, 299, 300]. RL, a key training method, introduces risks such as reward hacking, generalization failures, and language mixing, which can lead to harmful outcomes. Ensuring the safety of such systems like DeepSeek-R1 is urgent. While RL enhances reasoning, its uncontrollable nature raises concerns about safely guiding these models. SFT addresses some issues but is not a complete solution. A hybrid approach combining RL and SFT is needed to reduce harmful outputs while maintaining model effectiveness [301].